# IntroStyle: Training-Free Introspective Style Attribution using Diffusion Features

### v1

@@ Abstract
Text-to-image (T2I) models have gained widespread adoption among content creators and the general public. However, this has sparked significant concerns regarding data privacy and copyright infringement among artists. Consequently, there is an increasing demand for T2I models to incorporate mechanisms that prevent the generation of specific artistic styles, thereby safeguarding intellectual property rights. Existing methods for style extraction typically necessitate the collection of custom datasets and the training of specialized models. This, however, is resource-intensive, time-consuming, and often impractical for real-time applications. Moreover, it may not adequately address the dynamic nature of artistic styles and the rapidly evolving landscape of digital art. We present a novel, training-free framework to solve the style attribution problem, using the features produced by a diffusion model alone, without any external modules or retraining. This is denoted as introspective style attribution (IntroStyle) and demonstrates superior performance to state-of-the-art models for style retrieval. We also introduce a synthetic dataset of Style Hacks (SHacks) to isolate artistic style and evaluate fine-grained style attribution performance.
Code: https://github.com/AnandK27/introstyle
@@ 1 Introduction
Figure 1: Introspective Style Attribution (IntroStyle). Existing SOTA methods, e.g., CSD, are biased to content semantics and fail to retrieve images of similar styles, underscoring the need for a better style similarity metric. The top two rows show examples on the WikiArt [33] dataset and our proposed synthetic Style Hacks (SHacks) dataset, respectively, where a reference image, top-3 retrieval results of our method, top-2 retrieval results of CSD are shown from left to right. More importantly, our proposed method can be used as a metric for style measurement, as shown in the third row, with a lower score indicating images further away in style from the reference in the first column. Green colors indicate correct and red for incorrect retrievals.
@@
Diffusion models have significantly advanced image synthesis by applying iterative denoising processes guided by input prompts. Models like Stable Diffusion [30], DALL-E 3 [1], and Imagen [32] have emerged as a powerful paradigm for text-to-image synthesis, demonstrating remarkable capabilities in generating high-quality images from textual descriptions. The success of diffusion-based approaches has led to their application in various domains, including layout-to-image generation, text-guided image generation, and even video synthesis. However, these models’ widespread adoption and impressive performance have also raised concerns about problems like copyright infringement. For effective performance, the models require large-scale pre-training on diverse datasets [34]. Since these datasets are collected automatically and dominantly from the web, it is difficult to control image provenience and avoid the collection of copyrighted imagery. This problem is compounded by the tendency of diffusion models to replicate elements from their training data [2]. The legal implications of training models on copyrighted images have become a subject of recent debate and litigation, with artists arguing that the unauthorized use of their works for AI training [7, 12, 10] constitutes copyright infringement [17, 29].
@@
Current mitigation approaches include “unlearning” techniques to remove specific styles from AI models [44], though these require costly model retraining and may not fully address indirect style replication through alternative prompts [24]. Style “cloaking” methods [35] offer artists some protection but compromise the authentic viewing experience and place the burden of protection on creators. An alternative solution is to rely on attribution techniques [43, 37, 44], which can be used to attribute synthesized images to different artistic styles, allowing the post-hoc assessment of how much they might, for example, replicate an artist’s artwork. These methods are also more practical since they simply return examples from the dataset used to train the model, ranked by similarity to the synthesized image.
@@
Prior attribution approaches have proposed retraining the diffusion model [44] or adding external models to evaluate data similarity [43, 37]. This can add significant complexity since effective similarity models, like CLIP [26] or DINO [3], can be comparable in size to the diffusion model. On the other hand, prior diffusion model research has shown that different diffusion model layers develop distinct representations of higher and lower level image properties, such as structure vs color [42], producing features that can be leveraged to solve problems like image correspondence [39] or segmentation [40] without retraining. This led us to hypothesize that it may be possible to solve the attribution problem using the features produced by a diffusion model alone, without any external modules or retraining. We refer to this as introspective style attribution (IntroStyle).
@@
In this work, we show that the hypothesis holds. We propose a simple and effective approach to differentiate image styles by analyzing the similarity of the features produced by a diffusion model for different images. We frame the denoising network of the diffusion model as an autoencoder whose encoder produces stylistic features. We then propose a metric to compare these features that, while quite simple, are shown to be quite effective for style retrieval. In fact, a comparison with the current state-of-the-art on popular style retrieval datasets shows that IntroStyle significantly outperforms prior models trained in the existing evaluation datasets, as illustrated in the top rows of Fig. 1. Though the existing models are fine-tuned for style similarity, they still have semantic bias, whereas our method focuses solely on style.
@@
Beyond models, the problem of style attribution suffers from a shortage of datasets for evaluation. Creating such datasets is far from trivial since even the concept of “style” in art remains debatable. Nevertheless, many recognized artistic styles, such as surrealism and impressionism, are often associated with specific artists or movements. We use this social construct to define style as the collective global characteristics of an image identifiable with a particular artist or artistic movement. These characteristics encompass various elements such as color usage, brushstroke techniques, composition, and perspective. Under this definition, a single artist can produce art in several styles. Current datasets inadequately address AI-generated copyright concerns, primarily focusing on comparisons between authentic artworks rather than evaluating against AI-synthesized images. While CSD [37] introduces synthetic data for generated image evaluation and GDA[43] uses custom diffusion to generate training datasets, these existing datasets fail to account for prompt manipulation techniques that can circumvent copyright protections [41, 4, 45].
To address these problems, we create the Style Hacks (SHacks) dataset with inspiration from [4]. This dataset is based on real image queries that rank synthetic images created by manipulating their prompts. For each image, we produce a set of “innocent” prompts, which do not mention style, and a set of “hacked” prompts, which indirectly mention style in a way that is not easy to detect by text analysis. These prompts are then fed to a diffusion model to create corresponding images. By measuring similarities between features extracted from the query and synthetic images, it is then possible to perform fine-grained evaluations of the suitability of the feature representation for style attribution. Hence, style retrieval experiments on this dataset help evaluate the capability of different models to address problems like copyright infringement.
The second row of Fig. 1 shows retrieval results on the Style Hacks (SHacks) dataset, where the proposed model outperforms a state-of-the-art approach. Furthermore, the proposed style attribution approach allows improved sorting of images by similarity to a reference. An example is shown in the last row of Fig. 1, showing how images of style closer to the reference have higher similarity scores. This can be leveraged for style-based rejection to prevent diffusion models from generating images similar to a specific style.
@@
    We introduce the SHacks dataset to isolate artistic style and evaluate fine-grained style attribution performance.
@@ 2.1 Style Transfer
Style attribution and retrieval are related to style transfer, where the goal is to render one image in the style of another. Early approaches include [49], which formulated style transfer using Markov random fields. However, the topic raised in popularity with the rise of deep learning and the work of [9], which introduced Gram Matrices as style descriptors and an optimization framework for style transfer. Many variants have been proposed, including various regularizations to prevent distortions in reconstructed images [20] or representations of correlation beyond Gram matrices [5]. However, recent attribution work shows that Gram-style representations cannot guarantee effective style attribution [37].
Our work is instead inspired by more recent advances in style transfer using Generative Adversarial Networks (GANs) and other generative models, where methods like Adaptive Instance Normalization (AdaIN), which aligns the mean and variance of content features with those of style features, have been shown beneficial for real-time style transfer [14]. These ideas are central to the StyleGAN [15], where AdaIN layers are embedded into the generator architecture. This allows fine-grained control over different aspects of the generated images, from coarse features like pose to fine details like color schemes.
Recent advances in diffusion-based style transfer demonstrate promising results, notably through AdaIN layer manipulation of initial timestep latent [6], and DDIM Inversion-based style latent reuse [21]. These approaches informed our investigation into diffusion model latent distributions.
@@ 2.3 Data Attribution
While there is extensive literature on image retrieval, most methods focus on semantic similarity, i.e., the retrieval of images of similar content in terms of objects, scenes, etc. Style is a subtle property not well captured by representations, such as DINO [3] or CLIP [26], developed for this purpose. Few techniques have been developed specifically for style retrieval or attribution. Lee et al. [18] employed two separate neural network modules to facilitate style retrieval for image style and content. Recent works [31, 37, 43] instead train attribution models using contrastive objectives based on either Behance Artistic Media (BAM) dataset [31], a subset LAION Aesthetic Dataset [34] or synthetic style pairs. [31] proposes a dual encoder to separate content and style with style features obtained using AdaIN layers, whereas CSD[37] and GDA[43] fine-tune a pre-trained CLIP or DINO backbone with a linear layer on top with the former training the entire network and latter training the linear layer alone. It is also possible to formulate data attribution as the unlearning of target images [44], but this requires a different model for each target. Unlike these methods, we seek an introspective solution to style attribution that does not require external models or attribution-specific training.
@@ 2.4 Diffusion Features
In the diffusion model literature, various works have studied the representations learned for denoising, showing that different layers of the network learn features that capture different image properties, like structure, color, or texture [42, 50, 36]. Denoising Diffusion Autoencoders (DDAE) [46] leverage intermediate activations from pre-trained diffusion models for downstream tasks, demonstrating effectiveness in image classification through linear probing and fine-tuning. Recent advances include REPresentation Alignment (REPA) [47], which enhances semantic feature learning by aligning diffusion transformer representations with self-supervised visual features. Additionally, diffusion model features have shown promise in zero-shot applications like correspondence [39] and segmentation [40]. These observations have inspired our hypothesis that style attribution can be solved introspectively.
@@ 3 Method
Figure 2: IntroStyle computation of style similarity: channel-wise mean $\mu$ and variance $\sigma^{2}$ are computed for the identified style layers. Then a distance metric, 2-Wassertein Distance, can be used to measure styles between a pair of images.
@@ 3.2 Style attribution
Style attribution aims to identify in a reference dataset ${\cal D}=\{x_{1},\ldots,x_{n}\}$ of images $x_{i}$, usually but not necessarily the dataset used to train the diffusion model, the subset of images of style closest to that of a query image $q$, which is usually an image synthesized by the model. Like all retrieval operations, this requires two components: 1) a feature representation ${\cal F}:x\rightarrow f$ that maps each image $x$ into a feature tensor $f$ that captures the stylistic elements of $x$, and 2) a similarity metric $d({\cal F}(q),{\cal F}(x))$ that quantifies the style similarity of the feature tensors extracted from query $q$ and reference $x$ images. This allows retrieving the images in $\cal D$ that are most similar to $q$, which can then be inspected, e.g. to determine copyright violations.
@@ 3.3 Introspective style attribution
Style attribution is not a widely studied problem in the literature. Previous approaches [43, 37] attempted to train an external feature representation $\cal F$ or adapt existing feature representations, such as DINO [26], by addition of parameters specific to the extraction of style features. This has two major drawbacks: it requires computationally intensive large-scale training to cover the diffusion model’s sampling distribution, and the external feature extractor $\cal F$ introduces substantial memory and computational overhead—often exceeding the diffusion model’s resource requirements.
@@
Figure 3: Features from the first up-block are used to extract style features. Stable Diffusion 2.1 is used in our experiments.
a.1 Pablo Picasso’s Girl before a Mirror
a.2 Create a Cubist-style portrait of a woman facing her reflection in a mirror. Use bold, contrasting colors and geometric shapes to depict both the figure and her distorted image.
a.3 Create a portrait of a woman facing her reflection in a mirror.
b.1 Claude Monet’s Impression, Sunrise
b.2 Create an Impressionist-style seascape at dawn, featuring a hazy orange sun reflecting on rippling water. Include silhouettes of boats and misty harbor structures in muted blues and grays.
b.3 Create a seascape at dawn.
Figure 5: Style Hacks (SHacks) dataset samples. Each row shows images generated for a reference image by an artist. Reference image, three images generated by “hacked” prompts, and three images generated by “innocent” prompts are shown from left to right.
@@
To compute a feature representation, image $I$ is first mapped into a latent vector $z_{0}=\mathcal{E}(I)$, which is noised to a time $t$ using Eq. 1. The noised latent $z_{t}$ is then passed through the denoising network $\epsilon_{\theta}$ with null text embedding $c=\varnothing$. A feature tensor $F^{t}$ is then extracted from intermediate network layers. Given the UNet architecture commonly used to implement $\epsilon_{\theta}$, we consider this model an autoencoder that stacks a sequence of feature encoder and decoder blocks. The encoder/decoder split is quantified by equating all layers up to the upsampling block $idx$ as an encoder and the remaining upsampling layers as a decoder. The feature tensor $F^{t,idx}$ is finally extracted from block $idx$, as illustrated in Fig. 3 for $idx=1$.
@@
where $C$ is the number of channels at block $idx$. Both $t$ and $idx$ are hyperparameters of the method, which is summarized in Figures 2 and 3.
@@
The IntroStyle representation of an image models its feature statistics with a multivariate Gaussian distribution of $C$ dimensions. As shown in Figure 2, Gaussian parameters $(\mu_{1},\Sigma_{1})$ and $(\mu_{2},\Sigma_{2})$ are extracted for images $I_{1}$ and $I_{2}$, respectively. Hence, any similarity measure between two multivariate Gaussians can be used to compare the images. We have investigated several similarity measures and found the 2-Wasserstein ($W_{2}$) distance
@@
to be effective for style retrieval. A comparison to other metrics, including Euclidean ($L_{2}$) Distance, Gram Matrix and Jensen-Shannon Divergence ($JSD$) is presented in the experiments. In all cases, the covariance matrices are assumed diagonal with the variances of the IntroStyle representation as diagonal entries.
@@
Figure 6: Image Retrieval on WikiArt Dataset for IntroStyle, CSD, and GDA. The first column (left to right) shows the query image, followed by IntroStyle’s top 3 retrieval results, CSD and GDA, respectively. The colored boxes convey retrieval groundtruth similar to Fig. 1
@@
Table 1: Evaluation Results on WikiArt and DomainNet datasets. IntroStyle outperforms all previous methods significantly. The best values are highlighted with red and second best with orange.
4 The SHacks Dataset
To address limitations in existing attribution datasets, we introduce the Style Hacks (SHacks) dataset for evaluating style hacking detection. We selected prompt-image pairs for $25$ prominent artists from the LAION Aesthetic Dataset [34], focusing on their most recognized works as queries. We used Stable Diffusion v2.1 for each artist to synthesize a reference dataset containing images produced in response to two types of prompts:“innocent” and “hacked”. A set of such prompts was first generated for each prompt in the query set. “Innocent” prompts solely include a description of the image in terms of objects, scenes, etc., without mention of any style. “Hacked” prompts indicate the artist’s style indirectly, without explicitly referencing an artist name or painting [^1] It is assumed that if the user directly mentions the artist, a simple text filter will suffice to prevent generation.. Both “hacked” and “innocent” prompts were generated from the original prompts using ChatGPT[22] and manually inspected to remove any mention of the artist or their paintings directly. We generated $3$ prompts of each of the two types per original prompt and used the diffusion model to generate $12$ images per prompt. This produced a reference dataset with a total of $1800$ images. Samples of images generated using this method are shown in Fig. 5.
The reference images are ranked by similarity to each query image to evaluate style attribution performance. Ideally, this would rank all the “hacked” prompt images above the “innocent” ones. Hence, a success is declared when a “hacked” image is retrieved. Standard precision and recall metrics then evaluate style attribution performance.

### v2

@@ Abstract
Text-to-image (T2I) models have recently gained widespread adoption. This has spurred concerns about safeguarding intellectual property rights and an increasing demand for mechanisms that prevent the generation of specific artistic styles. Existing methods for style extraction typically necessitate the collection of custom datasets and the training of specialized models. This, however, is resource-intensive, time-consuming, and often impractical for real-time applications. We present a novel, training-free framework to solve the style attribution problem, using the features produced by a diffusion model alone, without any external modules or retraining. This is denoted as Introspective Style attribution (IntroStyle) and is shown to have superior performance to state-of-the-art models for style attribution. We also introduce a synthetic dataset of Artistic Style Split (ArtSplit) to isolate artistic style and evaluate fine-grained style attribution performance. Our experimental results on WikiArt and DomainNet datasets show that IntroStyle is robust to the dynamic nature of artistic styles, outperforming existing methods by a wide margin. Website: https://anandk27.github.io/IntroStyle
@@ 1 Introduction
Figure 1: Introspective Style Attribution (IntroStyle). Top two rows: comparison of style attribution by IntroStyle and a SOTA method (CSD). A query image, the top-3 retrieval results of IntroStyle, and the top-2 retrieval results of CSD are shown from left to right. Green colors indicate correct and red incorrect retrievals. First row: WikiArt [34] dataset. Second row: proposed synthetic Artistic Style Split (ArtSplit) dataset. While the SOTA style attributions are biased by image semantics, failing to retrieve images on the style of the query, IntroStyle retrieves images of the correct style. Bottom: retrieval scores of IntroStyle for ArtSplit synthetic images generated to have the semantic of the query (“Starry Night”) but different styles, demonstrating the effectiveness of IntroStyle as a metric for style measurement.
@@
Diffusion models have significantly advanced image synthesis by applying iterative denoising processes guided by input prompts. Models like Stable Diffusion [31], DALL-E 3 [1], and Imagen [33] have emerged as a powerful paradigm for T2I synthesis, demonstrating remarkable capabilities in generating high-quality images from textual descriptions. The success of diffusion-based approaches has led to their application in various domains, including layout-to-image generation, text-guided image generation, and even video synthesis. However, the widespread adoption and impressive performance of T2I modelss have also raised concerns about problems like copyright infringement. For effective performance, the models require large-scale pre-training on diverse datasets [35]. Since these datasets are collected automatically and dominantly from the web, it is difficult to control image provenance and avoid the collection of copyrighted imagery. This problem is compounded by the tendency of diffusion models to replicate elements from their training data [2]. The legal implications of training models on copyrighted images have become a subject of recent debate and litigation, with artists arguing that the unauthorized use of their works for AI training [6, 11, 9] constitutes copyright infringement [15, 30].
@@
Current mitigation approaches include “unlearning” techniques to remove specific styles from AI models [44], though these require costly model retraining and may not fully address indirect style replication through alternative prompts [25]. Style “cloaking” methods [36] offer artists some protection but compromise the authentic viewing experience and place the burden of protection on creators. An alternative solution is to rely on attribution techniques [43, 38, 44], which can be used to attribute synthesized images to different artistic styles, allowing the post-hoc assessment of how much they might, for example, replicate an artist’s artwork. These methods are also more practical since they simply return examples from the dataset used to train the model, ranked by similarity to the synthesized image. The main difficulty is to develop similarity functions that focus on style and are not confused by other image properties, such as similar image content or semantics.
@@
To solve this problem, prior attribution approaches have proposed retraining the diffusion model [44] or adding external models to evaluate data similarity [43, 38]. This can add significant complexity since effective similarity models, like CLIP [27] or DINO [3], can be comparable in size to the diffusion model. On the other hand, prior diffusion model research has shown that different diffusion model layers develop distinct representations of higher and lower level image properties, such as structure vs color [42], producing features that can be leveraged to solve problems like image correspondence [40] or segmentation [41] without retraining. This led us to hypothesize that it may be possible to solve the attribution problem using the features produced by a diffusion model alone, without any external modules or retraining. We refer to this as Introspective Style attribution (IntroStyle).
@@
In this work, we show that the hypothesis holds. We propose a simple and effective approach to differentiate image styles by analyzing the similarity of the features produced by a diffusion model for different images. We frame the denoising network of the diffusion model as an autoencoder whose encoder produces stylistic features. We then propose a metric to compare these features. While quite simple, this is shown to be quite effective for style retrieval. In fact, a comparison with the current state-of-the-art on popular style retrieval datasets shows that IntroStyle significantly outperforms prior models trained in the existing evaluation datasets, as illustrated in the top rows of Fig. 1. Although fine-tuned for style similarity, existing models still tend to retrieve images of similar semantics, whereas IntroStyle focuses much more uniquely on style.
@@
Beyond models, the problem of style attribution suffers from a shortage of evaluation datasets. Creating such datasets is far from trivial since even the concept of “style” in art remains debatable. Nevertheless, many recognized artistic styles, such as surrealism and impressionism, are often associated with specific artists or movements. We use this social construct to define style as the collective global characteristics of an image identifiable with a particular artist or artistic movement. These characteristics encompass various elements such as color usage, brushstroke techniques, composition, and perspective. Under this definition, a single artist can produce art in several styles. Current datasets, primarily focusing on comparisons between authentic artworks, provide limited ability for evaluating how strongly style attribution methods disentangles style from semantics. From this point of view, an ideal dataset for evaluating style attributions should contain images with all pairs of styles and semantics, a requirement only attainable by synthetic data. While CSD [38] introduces synthetic data for generated image evaluation and GDA [43] uses custom diffusion to generate training datasets, these datasets lack fine-grained separation of artistic style and image semantics, which are heavily correlated in their images. This makes the resulting models poorly suited for enforcing copyright protections [17].
To address these problems, we introduce the Artistic Style Split (ArtSplit) dataset, inspired by [38]. This dataset combines real artistic image queries with synthetic retrieval images generated via diffusion models using carefully crafted prompts. For each authentic painting, we create two distinct prompts: a “semantic” prompt that excludes stylistic elements and a “style” prompt that omits semantic content. These prompts are fed into a diffusion model to generate corresponding images, enabling comprehensive coverage of the style $\times$ semantics space encompassing all combinations in real artistic queries. By analyzing the similarity between features extracted from query and synthetic images, we can precisely evaluate how feature representations balance style attribution against semantic similarity. Hence, style retrieval experiments on this dataset provide clear insights into how different models disentangle style from semantics. This is illustrated in the last row of Fig. 1, where the painting Starry Night by Vincent Van Gogh is used to retrieve images of paintings of the Starry Night semantic under various styles, using IntroStyle. Note how images of style closer to the reference have higher similarity scores, providing evidence that the IntroStyle representation favors style over semantics. This can be leveraged to implement style-based rejection sampling, preventing diffusion models from generating images of a specific style.
@@
    We introduce the ArtSplit dataset to isolate artistic style and evaluate fine-grained style attribution performance.
@@ 2.1 Style Transfer
Style attribution and retrieval are related to style transfer, where the goal is to render one image in the style of another. While initially formulated using Markov random fields [48], style transfer raised in popularity with the rise of deep learning and the work of [8], which introduced Gram Matrices as style descriptors and an optimization framework for style transfer. Many variants have been proposed, including various regularizations to prevent distortions in reconstructed images [19] or representations of correlation beyond Gram matrices [4]. However, recent attribution work shows that Gram-style representations cannot guarantee effective style attribution [38].
Our work is inspired by more recent advances in style transfer using Generative Adversarial Networks (GANs), where methods like Adaptive Instance Normalization (AdaIN), which aligns the mean and variance of content features with those of style features, have been shown beneficial for real-time style transfer [12]. These ideas are central to the StyleGAN [13], where AdaIN layers are embedded into the generator architecture. This allows fine-grained control over different aspects of the generated images, from coarse features like pose to fine details like color schemes. More recently, diffusion-based style transfer has been implemented through AdaIN layer manipulation of initial timestep latent [5], and DDIM Inversion-based style latent reuse [20]. These approaches informed our investigation into diffusion model latent distributions.
@@ 2.3 Data Attribution
While there is extensive literature on image retrieval, most methods focus on semantic similarity, i.e., the retrieval of images of similar content in terms of objects, scenes, etc. Style is a subtle property not well captured by representations, such as DINO [3] or CLIP [27], developed for this purpose. Few techniques have been developed specifically for style retrieval or attribution. Lee et al. [16] employed two separate neural network modules to facilitate style retrieval for image style and content. Recent works [32, 38, 43] instead train attribution models using contrastive objectives based on either Behance Artistic Media (BAM) dataset [32], a subset LAION Aesthetic Dataset [35] or synthetic style pairs. Ruta et al. [32] proposes a dual encoder to separate content and style with style features obtained using AdaIN layers, whereas CSD [38] and GDA [43] fine-tune a pre-trained CLIP or DINO backbone with a linear layer on top with the former training the entire network and latter training the linear layer alone. It is also possible to formulate data attribution as the unlearning of target images [44], but this requires a different model for each target. Unlike these methods, we seek an introspective solution to style attribution that does not require external models or attribution-specific training.
@@ 2.4 Diffusion Features
In the diffusion model literature, various works have studied the representations learned for denoising, showing that different layers of the network learn features that capture different image properties, like structure, color, or texture [42, 49, 37, 22]. Denoising Diffusion Autoencoders (DDAE) [45] leverage intermediate activations from pre-trained diffusion models for downstream tasks, demonstrating effectiveness in image classification through linear probing and fine-tuning. Recent advances include REPresentation Alignment (REPA) [46], which enhances semantic feature learning by aligning diffusion transformer representations with self-supervised visual features. Additionally, diffusion model features have shown promise in zero-shot applications like correspondence [40], segmentation [41] and classification [21]. These observations have inspired our hypothesis that style attribution can be solved introspectively.
@@ 3 Method
Figure 2: IntroStyle computation of style similarity: channel-wise mean $\mu$ and variance $\sigma^{2}$ are computed for the identified style layer features. Then 2-Wasserstein Distance is used to measure style similarities between a pair of images.
@@ 3.2 Style attribution
Style attribution aims to identify in a reference dataset ${\cal D}=\{x_{1},\ldots,x_{n}\}$ of images $x_{i}$, usually but not necessarily the dataset used to train the diffusion model, the subset of images of style closest to that of a query image $q$, which is usually an image synthesized by the model. Like all retrieval operations, this requires two components: 1) a feature representation ${\cal F}:x\rightarrow f$ that maps each image $x$ into a feature tensor $f$ that captures the stylistic elements of $x$, and 2) a similarity metric $d({\cal F}(q),{\cal F}(x))$ that quantifies the style similarity of the feature tensors extracted from query $q$ and reference $x$ images. This allows retrieving the images in $\cal D$ that are most similar to $q$, which can then be inspected, e.g., to determine copyright violations.
@@ 3.3 Introspective style attribution
[…]
@@
Originals Synthesized
Semantic
Claude Monet’s Pablo Picasso’s Gustav Klimt’s Van Gogh’s Leonid Afremov’s
Impression, Sunrise Les Demoiselles Portrait of Adele Starry Night Melody Of The Night
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Claude Monet’s Impression, Sunrise Style: Claude Monet
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Pablo Picasso’s Les Demoiselles Style: Pablo Picasso
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Gustav Klimt’s Portrait of Adele Bloch-Bauer I Style: Gustav Klimt
Figure 3: Artistic Style Split (ArtSplit) dataset samples. Each row shows images generated with the same style, and each column with the same semantics.
@@
To compute a feature representation, image $I$ is first mapped into a latent vector $z_{0}=\mathcal{E}(I)$, which is noised to a time $t$ using Eq. 1. The noised latent $z_{t}$ is then passed through the denoising network $\epsilon_{\theta}$ with null text embedding $c=\varnothing$. A feature tensor $F^{t}$ is then extracted from intermediate network layers. Given the UNet architecture commonly used to implement $\epsilon_{\theta}$, we consider this model an autoencoder that stacks a sequence of feature encoder and decoder blocks. The encoder/decoder split is quantified by equating all layers up to the upsampling block idxidx as an encoder and the remaining upsampling layers as a decoder. The feature tensor $F^{t,idx}$ is finally extracted from block idxidx, as illustrated in Fig. 2 for $idx=1$.
@@
where $C$ is the number of channels at block idxidx. Both $t$ and idxidx are hyperparameters of the method, which is summarized in Fig. 2.
@@
The IntroStyle representation of an image models its feature statistics with a multivariate Gaussian distribution of $C$ dimensions. As shown in Fig. 2, Gaussian parameters $(\mu_{1},\Sigma_{1})$ and $(\mu_{2},\Sigma_{2})$ are extracted for images $I_{1}$ and $I_{2}$, respectively. Hence, any similarity measure between two multivariate Gaussians can be used to compare the images. We have investigated several similarity measures and found the 2-Wasserstein ($W_{2}$) distance
@@
to be effective for style retrieval. A comparison to other metrics, including Euclidean ($L_{2}$) Distance, Gram Matrix, and Jensen-Shannon Divergence ($JSD$), is presented in the experiments. In all cases, the covariance matrices are assumed diagonal with the variances of the IntroStyle representation as diagonal entries.
@@
Figure 4: Image retrieval on WikiArt Dataset for IntroStyle, CSD, and GDA. The first column (left to right) shows the query image, followed by IntroStyle’s top 3 retrieval results, CSD and GDA, respectively. Green colors indicate correct and red for incorrect retrievals.
@@
Table 1: Evaluation on WikiArt and DomainNet datasets. IntroStyle outperforms all previous methods significantly. The best values are highlighted with red and second best with orange.
4 The ArtSplit Dataset
To address the limitations of existing datasets for fine-grained evaluation of style retrieval, we propose the Artistic Style Split (ArtSplit) dataset. This was created with the prompt-image pairs of the $2$ most recognized works of $50$ prominent artists from the LAION Aesthetic Dataset [35]. For each of the two paintings, ChatGPT[23] was asked to generate a “style” specification and a “semantic” description, such that there is no style information in the semantic description and vice-versa. Stable Diffusion v2.1 was then used with a combination of two prompts, “style” and “semantic,” to synthesize a reference image dataset. With 50 artists (“styles”) and 100 paintings (“semantics”), this led to $50\times 100=5,000$ prompt combinations. A set of $12$ images was sampled per combination, yielding $60,000$ images in total. The procedure is detailed in Supplemental Section D.
Fig. 3 shows some of the images in the dataset. The left column shows three of the $100$ original paintings used to produce the style and semantic descriptions, e.g. style: “Gustav Klimt” and semantic: “Portrait of Adele”. The remaining table entries are images synthesized for different combinations of style (row) and semantic (column). Note how the procedure creates a fine-grained sample of the style $\times$ semantic space.
Style attribution performance is measured using the original artist’s paintings (left column at Fig. 3) to retrieve images. Each original is used as a query, the dataset images are ranked by similarity to this query, and performance is evaluated by the agreement between the labels of the retrieved images and that of the query. This is measured using a weighted score of Top-k retrieved neighbors that are correctly matched. Based on the labels, we perform two assessments on the retrieved images. Style labels allow the evaluation of style attribution accuracy. On the other hand, semantic labels allow the evaluation of semantic similarity and reveal the model’s semantic bias. A good representation for style attribution should simultaneously achieve high style and low semantic retrieval accuracy.